# A Practical Protocol for Quantum Oblivious Transfer from One-Way Functions

### v1

@@ 5 OT from equivocal and relaxed-extractable bit commitment
In this section, we build an oblivious transfer functionality assuming an equivocal and ($\chi$, $\epsilon$)-relaxed extractable bit commitment, with $\chi=\omega(\log^{2}(k)/k)$. The structure is similar to [DFL+09, BCKM21], but applying it to our setting. Lastly, we show how to optimize the number of runs of OT in Corollary 2.
@@
4:$A$ and $B$ execute ERE-Decommitment sequentially for every $i\in T$. Then, $A$ obtains $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in T}$.
@@ 5.1 Security against a malicious Alice
7:$\text{Sim}_{A^{*}}$ obtains $(s_{0},PRG(h(s_{0},\boldsymbol{x}_{0}))\oplus m_{0},s_{1},PRG(h(s_{1},\boldsymbol{x}_{1}))\oplus m_{1})$, and computes $m_{0}$ and $m_{1}$.
@@
It follows from the fact that the delay in the measurement of $\psi_{i}$ with $i\in T$ cannot be noticed by $A^{*}$ since, in both cases, the measurement bases are sampled at random, similar to Lemma 5. ∎
@@ 5.2 Security against a malicious Bob
Consider a QOT protocol between an honest Alice and a malicious Bob (Algorithm 5) with ERE-Commitment as the bit commitment subprotocol (Algorithm 3). Let $b\in\{0,1\}$ be a random bit and $K_{b}=h(r,\boldsymbol{x}_{I_{b}})\in\{0,1\}^{\ell}$ and $K_{\bar{b}}=h(r,\boldsymbol{x}_{I_{\bar{b}}})\in\{0,1\}^{\ell}$, the keys. $h(r,x):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ is a 2-universal hash function, and $I_{0}$ and $I_{1}$ are the two subsets in which $\bar{T}\subset[2\lambda_{OT}]$ the unchallenged commitments are divided. Then, given $K_{b}$, for any sampling errors $\xi,\delta>0$ and an experimental error $\alpha\geq 0$:
$\begin{split}&\Delta(\rho_{K_{\bar{b}}K_{b}E},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{K_{b}E})\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\chi))\lambda_{OT}/2-\ell-q]}\\ &+\sqrt{6}\exp(-\lambda_{OT}\delta^{2}/100)+2\exp(-\xi^{2}\lambda_{OT}/2),\end{split}$ (5.3)
@@
After the commitment and check subprotocols, the next step in the QOT protocol is when Alice sends the bases $\boldsymbol{\theta}|_{\bar{T}}$ to Bob, and Bob partitions $\bar{T}$ into two sets, $I_{0}$ and $I_{1}$. There exists a bit $b\in\{0,1\}$ such that $d_{H}(\boldsymbol{\theta}|_{I{\bar{b}}},\boldsymbol{\hat{\theta}}|_{I{\bar{b}}})\geq(1/2-\epsilon)\lambda_{OT}/2$, regardless of how Bob splits the sets of indexes. By applying Equation 2.1 to our state $\rho_{X_{\bar{b}}A^{b}E}$, itself obtained from measuring the subsystem $A^{\bar{b}}$ on $\left|\psi_{A^{\bar{b}}A^{b}E}\right\rangle$, privacy amplification gives, for $\xi=2\gamma$:
$\begin{split}H_{\text{min}}(X_{\bar{b}}|A^{b}E)\geq&d_{H}(\boldsymbol{\theta}|_{I{\bar{b}}},\boldsymbol{\hat{\theta}}|_{I{\bar{b}}})-h_{2}(\delta+\alpha+\eta+\zeta)|I_{\bar{b}}|\\ &\geq\left(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\chi)\right)\frac{\lambda_{OT}}{2}.\end{split}$
Given a $(\chi,\epsilon)$-relaxed extractable ERE-Commitment, with $\chi=\eta+\zeta=\omega(\log^{2}(k)/k)$, ERE-Commitment is relaxed extractable with probability $\epsilon=1-negl(k)$. ∎
@@
1:$\text{Sim}_{B^{*}}$ chooses $\boldsymbol{x}\leftarrow\{0,1\}^{2\lambda_{OT}}$ and $\boldsymbol{\theta}\leftarrow\{+,\times\}^{2\lambda_{OT}}$ and sends $\left|\boldsymbol{x}\right\rangle_{\boldsymbol{\theta}}$ to $B$.
@@
4:$\text{Sim}_{B^{*}}$ and $B$ execute ERE-Decommitment sequentially for every $i\in T$. Then, $\text{Sim}_{B^{*}}$ obtains the set $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in T}$ and checks that $x_{i}=\hat{x}_{i}$ whenever $\theta_{i}=\hat{\theta}_{i}$, up to an experimental error $\alpha$. If the test does not pass, $\text{Sim}_{B^{*}}$ aborts.
@@
8:$\text{Sim}_{B^{*}}$ samples seeds $s_{0},s_{1}\leftarrow\{0,1\}^{k(\lambda_{PQS})}$ and sends $(s_{0},PRG(h(s_{0},\boldsymbol{x}_{0}))\oplus m_{0},s_{1},PRG(h(s_{1},\boldsymbol{x}_{1}))\oplus m_{1})$, where $\boldsymbol{x}_{0}=\{x_{i}:i\in I_{0}\}$ and $\boldsymbol{x}_{1}=\{x_{i}:i\in I_{1}\}$. $\text{Sim}_{B^{*}}$ outputs the final state of $R^{*}$.
@@
It follows from the $(\omega(\log^{2}(k)/k)),negl(k))$-relaxed-extractability definition of the bit commitment. ∎
@@
This would mean that there is a non-uniform quantum poly-time distinguisher $\mathcal{D}^{*}$ that can distinguish between the messages $m_{\bar{b}}$ obtained in $\text{Hyb}_{1}$ and $m_{\bar{b}}$ obtained in $\text{Hyb}_{2}$. Since $m_{\bar{b}}$ is encoded as $PRG(h(s_{\bar{b}},x_{\bar{b}}))\oplus m_{\bar{b}}$, this would be equivalent to distinguishing between the seeds $h(s_{\bar{b}},x_{\bar{b}})$ of $\text{Hyb}_{1}$ and $\text{Hyb}_{2}$. This leads to a contradiction since $h(s_{\bar{b}},x_{\bar{b}})$ is statistically close to a random string due to privacy amplification, as proven in Lemma 11. ∎
@@
$\begin{split}&\Delta(\rho_{K_{\bar{b}}K_{b}E},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{K_{b}E})\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\chi)v/2-\ell-q]}+\sqrt{6}\exp(-\lambda_{OT}\delta^{2}/100)+2\exp(-\xi^{2}\lambda_{OT}/2),\end{split}$
@@
$H_{\text{min}}(X_{\bar{b}}|A^{b}E)\geq(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\chi))\frac{v}{2}.$ (5.4)
@@ 6 Acknowledgements
ED, ABG and AY are supported by the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska Curie Grant No. 101072637, Project Quantum-Safe Internet (QSI). ED and PL acknowledge PEPR integrated project QCommTestbed ANR-22-PETQ-0011 part of Plan France 2030, VY acknowledges the support of European Research Council Starting Grant QUSCO, 758911. ED and AI acknowledge the support of QuantERA QuantaGENOMICS under Grant Agreement No. 101017733. ABG is supported by ANR JCJC TCS-NISQ ANR-22-CE47-0004. This work was done in part while ABG was visiting the Simons Institute for the Theory of Computing.
@@ Appendix 0.A Distance bound for a malicious Alice
Consider ERE-Commitment (see Algorithm 3), between an honest sender and a malicious receiver with a computationally hiding and ($\eta$, $2^{-\eta n}$)-relaxed statistically binding EqCommitment as base commitment. Let $K\coloneqq g(r,\boldsymbol{x})$ be the key in $\{0,1\}^{\ell}$ output by Alice, where $h(r,x):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ is a universal hash function, and $\bar{T}\subset[4\lambda_{EX}]$, the set of commitments that have not been challenged. Moreover, $m\in\bar{T}$ is a randomly chosen subset of non-challenged commitments that are input in the hash function. Then, for any sampling error $\xi,\delta>0$ and an experimental error $\alpha\geq 0$,
$\begin{split}&\Delta(\rho_{KE},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{E})\leq\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\eta))m-\ell-q]}+\sqrt{6}\exp(-2\lambda_{EX}\delta^{2}/100)+2\exp(-4\xi^{2}\lambda_{EX}),\end{split}$ (0.A.1)
@@
$H_{\text{min}}(X|E)\geq d_{H}(\boldsymbol{\theta}|_{\bar{T}},\boldsymbol{\hat{\theta}}|_{\bar{T}})-h_{2}(\delta+\alpha+\eta)|{\bar{T}}|\geq\left(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\eta)\right)2\lambda_{EX}.$ (0.A.2)
@@
$H_{\text{min}}(X_{m}|E)\geq\left(\frac{1}{2}-\xi-h_{2}(\delta+\alpha+\eta)\right)m.$ (0.A.3)
Given a $(\eta,2^{-k\eta})$-relaxed statistical binding EqCommitment, the probability of being relaxed statistical binding is $1-negl(k)$ when $\eta=\omega(\log^{2}(k)/k)$

### v2

@@ 5 OT from equivocal and relaxed-extractable bit commitment
In this section, we build an oblivious transfer functionality assuming an equivocal and $\chi$-relaxed extractable bit commitment, with $\chi=\omega(\log^{2}(k)/k)$. The structure is similar to [DFL+09, BCKM21], but applying it to our setting. Lastly, we show how to optimize the number of runs of OT in Corollary 2.
@@
4:$A$ and $B$ execute ERE-Decommitment for every $i\in T$. Then, $A$ obtains $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in T}$.
@@ 5.1 Security against a malicious Alice
7:$\text{Sim}_{A^{*}}$ obtains $(s_{0},PRG(h(s_{0},\boldsymbol{x}_{0}))\oplus m_{0},s_{1},PRG(h(s_{1},\boldsymbol{x}_{1}))\oplus m_{1})$, and computes $m_{0}$ and $m_{1}$. Both, $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$ are the corrected measured strings $\boldsymbol{\hat{x}}_{0}$ and $\boldsymbol{\hat{x}}_{1}$ with the syndromes $Synd_{0}$ and $Synd_{1}$.
@@
It follows from the fact that the delay in the measurement of $\psi_{i}$ with $i\in T$ cannot be noticed by $A^{*}$ since, in both cases, the measurement bases are sampled at random, similar to Lemma 6. Then, both extracted messages $m_{0}$ and $m_{1}$ are obtained by using the same error correction syndromes as in $Hyb_{1}$, leading to the same distribution. ∎
@@ 5.2 Security against a malicious Bob
Consider a QOT protocol between an honest Alice and a malicious Bob (Algorithm 5) with ERE-Commitment as the bit commitment subprotocol (Algorithm 3). Let $b\in\{0,1\}$ be a random bit and $K_{b}=h(r,\boldsymbol{x}_{I_{b}})\in\{0,1\}^{\ell}$ and $K_{\bar{b}}=h(r,\boldsymbol{x}_{I_{\bar{b}}})\in\{0,1\}^{\ell}$, the keys. $h(r,x):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ is a 2-universal hash function, and $I_{0}$ and $I_{1}$ are the two subsets in which $\bar{T}\subset[2\lambda_{OT}]$ the unchallenged commitments are divided. Then, given $K_{b}$, for any sampling errors $\xi,\delta>0$, a bit-flip error $\alpha\geq 0$ and a proportion $\vartheta\geq 0$ of leaked bits:
$\begin{split}&\Delta(\rho_{K_{\bar{b}}K_{b}E},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{K_{b}E})\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[\left(\frac{1}{2}-\xi-2\vartheta\right)\frac{\lambda_{OT}}{2}-h_{2}\left(\delta+\alpha+\chi\right)\frac{\lambda_{OT}}{2}\left(1-2\vartheta\right)-\ell-q]}\\ &+\sqrt{6}\exp(-\lambda_{OT}\delta^{2}/100)+2\exp(-\xi^{2}\lambda_{OT}/2),\end{split}$ (5.3)
@@
After the commitment and check subprotocols, the next step in the QOT protocol is when Alice sends the bases $\boldsymbol{\theta}|_{\bar{T}}$ to Bob, and Bob partitions $\bar{T}$ into two sets, $I_{0}$ and $I_{1}$. There exists a bit $b\in\{0,1\}$ such that $d_{H}(\boldsymbol{\theta}|_{I{\bar{b}}},\boldsymbol{\hat{\theta}}|_{I{\bar{b}}})\geq(1/2-\epsilon-2\vartheta)\lambda_{OT}/2$, regardless of how Bob splits the sets of indexes and taking the maximum advantage of the leaked bits, i.e., all the leaked bits are used in the set $I_{\bar{b}}$. Moreover, the real size of the non-sampled subset $I_{\bar{b}}$ for which the binary entropy in bounded is $\left(1-2\vartheta\right)$. By applying Equation 2.1 to our state $\rho_{X_{\bar{b}}A^{b}E}$, itself obtained from measuring the subsystem $A^{\bar{b}}$ on $\left|\psi_{A^{\bar{b}}A^{b}E}\right\rangle$, privacy amplification gives, for $\xi=2\gamma$:
$\begin{split}H_{\text{min}}(X_{\bar{b}}|A^{b}E)\geq&d_{H}(\boldsymbol{\theta}|_{I{\bar{b}}},\boldsymbol{\hat{\theta}}|_{I{\bar{b}}})-h_{2}(\delta+\alpha+\eta+\zeta)|I_{\bar{b}}|\\ &\geq\left(\frac{1}{2}-\xi-2\vartheta\right)\frac{\lambda_{OT}}{2}-h_{2}\left(\delta+\alpha+\chi\right)\frac{\lambda_{OT}}{2}\left(1-2\vartheta\right).\end{split}$
Given a $\chi$-relaxed extractable ERE-Commitment, with $\chi=\eta+\zeta=\omega(\log^{2}(k)/k)$, ERE-Commitment is relaxed extractable. ∎
@@
1:$\text{Sim}_{B^{*}}$ chooses $\boldsymbol{x}\leftarrow\{0,1\}^{2\lambda_{OT}}$ and $\boldsymbol{\theta}\leftarrow\{+,\times\}^{2\lambda_{OT}}$ and sends $\left|\boldsymbol{x}\right\rangle_{\boldsymbol{\theta}}$ to $B^{*}$.
@@
4:$\text{Sim}_{B^{*}}$ and $B^{*}$ execute ERE-Decommitment for every $i\in T$. Then, $\text{Sim}_{B^{*}}$ obtains the set $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in T}$ and checks that $x_{i}=\hat{x}_{i}$ whenever $\theta_{i}=\hat{\theta}_{i}$, up to an experimental error $\alpha$. If the test does not pass, $\text{Sim}_{B^{*}}$ aborts.
@@
8:$\text{Sim}_{B^{*}}$ samples seeds $s_{0},s_{1}\leftarrow\{0,1\}^{k(\lambda_{PQS})}$ and sends $(s_{0},PRG(h(s_{0},\boldsymbol{x}_{0}))\oplus m_{0},s_{1},PRG(h(s_{1},\boldsymbol{x}_{1}))\oplus m_{1})$, where $\boldsymbol{x}_{0}=\{x_{i}:i\in I_{0}\}$ and $\boldsymbol{x}_{1}=\{x_{i}:i\in I_{1}\}$. He also sends the syndromes $Synd_{0}$ and $Synd_{1}$ for correcting a proportion $\alpha$ of errors of the bit strings $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$. $\text{Sim}_{B^{*}}$ outputs the final state of $R^{*}$.
@@
It follows from the $\omega(\log^{2}(k)/k)$-relaxed-extractability definition of the bit commitment. ∎
@@
This would mean that there is a non-uniform quantum poly-time distinguisher $\mathcal{D}^{*}$ that can distinguish between the messages $m_{\bar{b}}$ obtained in $\text{Hyb}_{1}$ and $m_{\bar{b}}$ obtained in $\text{Hyb}_{2}$. Since $m_{\bar{b}}$ is encoded as $PRG(h(s_{\bar{b}},\hat{\boldsymbol{x}}_{\bar{b}}))\oplus m_{\bar{b}}$, this would be equivalent to distinguishing between the seeds $h(s_{\bar{b}},\hat{\boldsymbol{x}}_{\bar{b}})$ of $\text{Hyb}_{1}$ and $\text{Hyb}_{2}$. This leads to a contradiction since $h(s_{\bar{b}},\hat{\boldsymbol{x}}_{\bar{b}})$ is statistically close to a random string due to privacy amplification, as proven in Lemma 12. Moreover, in both distributions, the applied syndromes for the error correction of the strings $\hat{\boldsymbol{x}}_{\bar{b}}$ and $\hat{\boldsymbol{x}}_{b}$ are the same one, leading to the same corrected strings. ∎
@@
$\begin{split}&\Delta(\rho_{K_{\bar{b}}K_{b}E},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{K_{b}E})\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[(\frac{1}{2}-\xi)\frac{v}{2}-\vartheta\lambda_{OT}-h_{2}(\delta+\alpha+\chi)(\frac{v}{2}-\vartheta\lambda_{OT})-\ell-q]}\\ &+\sqrt{6}\exp(-\lambda_{OT}\delta^{2}/100)+2\exp(-\xi^{2}\lambda_{OT}/2),\end{split}$
@@
$H_{\text{min}}(X_{\bar{b}}|A^{b}E)\geq(\frac{1}{2}-\xi)\frac{v}{2}-\vartheta\lambda_{OT}-h_{2}(\delta+\alpha+\chi)(\frac{v}{2}-\vartheta\lambda_{OT}).$ (5.4)
@@ 6 Acknowledgements
AY acknowledges Lucas Hanouz for the discussions related with error correction. We also want to acknowledge Matías Bolaños for highlighting the necessity of adding multiphoton events and suggesting some improvements in the protocol. Lastly, we want to acknowledge anonymous reviewers for their feedback. ED, ABG and AY are supported by the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska Curie Grant No. 101072637, Project Quantum-Safe Internet (QSI). ED and PL acknowledge PEPR integrated project QCommTestbed ANR-22-PETQ-0011 part of Plan France 2030, VY acknowledges the support of European Research Council Starting Grant QUSCO, 758911. ED and AI acknowledge the support of QuantERA QuantaGENOMICS under Grant Agreement No. 101017733. ABG is supported by ANR JCJC TCS-NISQ ANR-22-CE47-0004. This work was done in part while ABG was visiting the Simons Institute for the Theory of Computing.
@@ Appendix 0.A Distance bound for a malicious Alice
Consider ERE-Commitment (see Algorithm 3), between an honest sender and a malicious receiver with a computationally hiding and ($\eta$, $2^{-\eta n}$)-relaxed statistically binding EqCommitment as base commitment. Let $K\coloneqq g(r,\boldsymbol{x})$ be the key in $\{0,1\}^{\ell}$ output by Alice, where $h(r,x):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ is a universal hash function, and $\bar{T}\subset[4\lambda_{EX}]$, the set of commitments that have not been challenged. Moreover, $m\in\bar{T}$ is a randomly chosen subset of non-challenged commitments that are input in the hash function. Then, for any sampling error $\xi,\delta>0$, a bit-flip error $\alpha\geq 0$ and a proportion $\vartheta\geq 0$ of leaked bits:
$\begin{split}&\Delta(\rho_{KE},\frac{1}{2^{\ell}}\text{I}\otimes\rho_{E})\leq\\ &\leq\frac{1}{2}\cdot 2^{-\frac{1}{2}[\left(\frac{1}{2}-\xi\right)m-2\vartheta\lambda_{EX}-h_{2}\left(\delta+\alpha+\eta\right)\left(m-2\vartheta\lambda_{EX}\right)-\ell-q]}\\ &+\sqrt{6}\exp(-2\lambda_{EX}\delta^{2}/100)+2\exp(-4\xi^{2}\lambda_{EX}),\end{split}$ (0.A.1)
@@
$\begin{split}H_{\text{min}}(X|E)&\geq d_{H}(\boldsymbol{\theta}|_{\bar{T}},\boldsymbol{\hat{\theta}}|_{\bar{T}})-h_{2}(\delta+\alpha+\eta)|{\bar{T}}|\\ &\geq\left(\frac{1}{2}-\xi-\vartheta\right)2\lambda_{EX}-h_{2}\left(\delta+\alpha+\eta)\right)2\lambda_{EX}(1-\vartheta).\end{split}$ (0.A.2)
@@
$H_{\text{min}}(X_{m}|E)\geq\left(\frac{1}{2}-\xi\right)m-2\vartheta\lambda_{EX}-h_{2}\left(\delta+\alpha+\eta\right)\left(m-2\vartheta\lambda_{EX}\right).$ (0.A.3)
Given a $\eta$-relaxed statistical binding EqCommitment, the probability of being relaxed statistical binding is $1-negl(k)$ when $\eta=\omega(\log^{2}(k)/k)$